# MathEDU: Towards Adaptive Feedback for Student Mathematical Problem-Solving

### v1

@@ 8 Conclusion
Assessment in mathematics education is critical. However, this grading process is often time-consuming and labor-intensive for teachers, particularly when it requires detailed analysis of each student’s individual problem-solving steps. This study constructed the MathEdu dataset, which includes real student problem-solving processes and teacher feedback for GRE-level math problems. We explore the use of LLMs to automate grading and provide adaptive feedback. While LLMs perform well in identifying correctness and errors, they struggle with generating personalized suggestion. An advanced method to enhance LLMs’ ability to understand and interpret mathematical reasoning is left as our future work.
@@
Limited Model and Method Exploration. Our study focused on Llama3 8B, Llama3 70B, and GPT-3.5, with experiments limited to few-shot prompting and LoRA-based fine-tuning. Other models, such as Gemma 2 (Team et al., 2024), Qwen 2.5 (Yang et al., 2024), and Phi-3 (Abdin et al., 2024), as well as newer versions of Llama, remain unexplored. Additionally, our fine-tuning methods did not delve into incorporating students’ individual learning progress, such as tracking their mastered concepts. Despite these limitations, we hope our work can contribute to the growing exploration of LLM applications in mathematics education and encourage further research to refine and expand these methods.
@@
To address potential privacy concerns associated with the MathEdu dataset, this section outlines the measures taken to protect annotators’ privacy and the ethical considerations in releasing the dataset. The MathEdu dataset contains problem-solving records annotated by students from diverse academic backgrounds. However, no personally identifiable information, such as names or IDs, was collected. Each record is assigned an anonymized identifier to categorize entries from the same student. While the dataset includes the students’ academic backgrounds (i.e., majors), this information is provided solely for analytical purposes. As highlighted in previous sections, understanding how problem-solving styles vary across mathematical abilities and expertise is a valuable area of research. All student annotators were informed that their personal data would remain confidential and that only their academic background would be disclosed. This consent was obtained prior to participation in the annotation task. The dataset’s structure ensures that individual privacy is preserved while facilitating research in educational applications. In alignment with ethical guidelines, the dataset will only be released for research purposes.
@@ B.1 Annotation Guidelines
Student Major Total General Gain Physics Geometry Probability Other
Student 1 Applied Mathematics 683 276 (40.41%) 132 (19.33%) 182 (26.64%) 49 (7.17%) 7 (1.02%) 37 (5.42%)
Student 2 Finance 685 297 (43.36%) 133 (19.42%) 188 (27.45%) 39 (5.69%) 4 (0.58%) 24 (3.50%)
Student 3 Japanese 678 281 (41.45%) 136 (20.06%) 180 (26.55%) 42 (6.19%) 7 (1.03%) 32 (4.72%)
Student 4 Information Management 660 268 (40.60%) 140 (21.21%) 167 (25.30%) 46 (6.97%) 8 (1.21%) 31 (4.70%)
Student 5 Mathematics Education 682 275 (40.32%) 139 (20.38%) 186 (27.27%) 41 (6.01%) 8 (1.17%) 33 (4.84%)
Student 6 Physics 660 262 (39.70%) 127 (19.24%) 175 (26.52%) 55 (8.33%) 7 (1.06%) 34 (5.15%)
Table 10: Student Majors and Distribution of Questions Answered.
Error Type Explanation
Wrong Mathematical Operation/Concept Student applies an incorrect mathematical operation or uses an inappropriate mathematical concept to solve a problem.
Calculation Error Mistakes in calculations, such as errors in solving equations, arithmetic mistakes, and incorrect unit conversions.
Incomplete Answer Student used a correct formula or procedure but did not complete it.
Careless Error Errors caused by students’ carelessness in answering, including number substitution errors and missing digits.
Lack of Necessary Mathematical Concepts Errors in answering caused by a lack of essential mathematical knowledge or techniques.
Table 11: Error Types and Explanations.
@@ B.2 Academic Backgrounds and Question Distribution of Student Annotators
To collect data on math problem-solving, we invited six university students to participate in the annotation process. Table 10 presents the majors of the student annotators and the distribution of questions they answered. We aimed for diversity by including individuals from different academic backgrounds, such as Japanese Studies and Applied Mathematics, to capture a range of math proficiency levels in the problem-solving process. All annotators involved in the task, including students and mathematics experts serving as teachers, were compensated based on the minimum hourly wage standards of their respective countries.
@@ Appendix C Supplementary Information for Each Error Category
Detailed descriptions of these error categories are provided in Table 11. The following are examples of each error category.
@@
Student Process: $46-36=10$ $10\,\mathrm{km/hr}\times\frac{5}{18}=\frac{100}{36}\,\mathrm{m/s}$ $\frac{100}{36}\times 144=400\,\mathrm{m}$
@@
Appendix D Details of Fine-Tuning Methods
We employ LoRA fine-tuning to train an LLM $\mathcal{M}$, parameterized by $\Phi$. The training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,\ldots,N}$. Given task-specific parameter increment $\Delta\Phi=\Delta\Phi(\Theta)$, we optimize over $\Theta$:
End-to-end without rationale: Similar to the “End-to-end with rationale” approach, this method also uses an end-to-end training approach but excludes the rationale in the input data. The training data used in LoRA fine-tuning is denoted as $\mathcal{Z}=\{(x_{i},y_{i})\}_{i=1,\ldots,N},$ where $x_{i}=\mathcal{P}_{D}(q_{i},p_{i})$ and $y_{i}=(c_{i},E_{i},T_{i})$. Here, $\mathcal{P}_{D}$ represents the prompt that takes $q_{i}$ and $p_{i}$ to generate grading results.
Appendix E Input Format
This section presents the few-shot prompts and fine-tuning prompts used for the answer accuracy assessment task. The prompt for rating the generated feedback is also included. Table 17 presents the few-shot prompts. Since we examined the models both with and without rationale, the prompts are presented in two formats. In these prompts, “few-shot examples” denotes the position for examples, where six samples are randomly selected from the training dataset, comprising three correct and three incorrect responses.
Without Rationale With Rationale
You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, identify where the student went wrong and provide explanations and advice. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. [wrong equation] ... [Teacher’s explanations and advice] ... {few-shot examples} [Question] : {question} Option: {options} [Student’s Answer] : {student process} You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, identify where the student went wrong and provide explanations and advice. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. [wrong equation] ... [Teacher’s explanations and advice] ... {few-shot examples} [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process}
Table 17: Few-Shot Prompts for Grading Student Problem-Solving Results.
Table 18 presents the end-to-end prompts. Since the fine-tuned models are trained and inferred using a zero-shot approach, there is no “few-shot examples” field included. The prompt content has been adjusted to meet the format required by Llama. Table 19 presents the prompts used for the single-task and multi-task models. Because we separated the tasks of answer accuracy assessment, problem-solving Error Identification, and feedback generation, we created three distinct prompts for each task.
Without Rationale With Rationale
### Instruction: You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, identify where the student went wrong and provide explanations and advice. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. [wrong equation] ... [Teacher’s explanations and advice] ... ### Input: [Question] : {question} Option: {options} [Student’s Answer] : {student process} ### Response: ### Instruction: You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, identify where the student went wrong and provide explanations and advice. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. [wrong equation] ... [Teacher’s explanations and advice] ... ### Input: [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process} ### Response:
Table 18: End-to-End Fine-Tuning Prompt for Grading Student Problem-Solving Results.
Category Without Rationale With Rationale
Answer Accuracy Assessment ### Instruction: You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] is correct or not. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. ### Input: [Question] : {question} Option: {options} [Student’s Answer] : {student process} ### Response: ### Instruction: You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] is correct or not. For correct answers, respond: The student’s answer is correct. For incorrect answers, respond: The student’s answer is incorrect. ### Input: [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process} ### Response:
Problem-Solving Error Identification ### Instruction: You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] contains any wrong equations. For correct answers, respond: The student’s answer doesn’t contain wrong equations. For incorrect answers with wrong equations, respond: [wrong equation] ... ### Input: [Question] : {question} Option: {options} [Student’s Answer] : {student process} ### Response: ### Instruction: You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] contains any wrong equations. For correct answers, respond: The student’s answer doesn’t contain wrong equations. For incorrect answers with wrong equations, respond: [wrong equation] ... ### Input: [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process} ### Response:
Feedback    Generation ### Instruction: You are a math teacher. According to the [Question], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, explain why the student went wrong and provide advice on how to correct their mistake. For correct answers, respond: The student’s answer is correct and doesn’t need advice. For incorrect answers, respond: [Teacher’s explanations and advice] ... ### Input: [Question] : {question} Option: {options} [Student’s Answer] : {student process} ### Response: ### Instruction: You are a math teacher. According to the [Question] and [Rationale], please indicate whether the [Student’s Answer] is correct or not. If the [Student’s Answer] is incorrect, explain why the student went wrong and provide advice on how to correct their mistake. For correct answers, respond: The student’s answer is correct and doesn’t need advice. For incorrect answers, respond: [Teacher’s explanations and advice] ... ### Input: [Question] : {question} Option: {options} [Rationale] : {rationale} [Student’s Answer] : {student process} ### Response:
Table 19: Single-Task and Multi-Task Learning Prompts for Grading Student Problem-Solving Results.
Table 20 presents the prompts used for rating the generated feedback, where we first outline the rating criteria, with scores ranging from 0 to 5, indicating poor to excellent. Additionally, we specify that the output should follow the format: “The rating for this correct suggestion is:” to facilitate the result collection process.
Given a problem, student’s process and the teacher’s manual suggestions, please rate the following feedback generated by different models on a scale from 0 to 5. The criteria for the rating are:
0: Completely irrelevant or nonsensical suggestions. 1: Very poor suggestions with little relevance or helpfulness. 2: Poor suggestions with some relevance but mostly unhelpful. 3: Average suggestions that are somewhat relevant and somewhat helpful. 4: Good suggestions that are relevant and helpful but could be improved. 5: Excellent suggestions that are highly relevant, helpful, and well-aligned with the teacher’s manual suggestions.
Please only answer in the following format:
The rating for this feedback is:
Table 20: Rating Prompt.
Train Validation Test
Time Series Split 2,836 609 603
Leave-One-Out S1 3,028 337 683
S2 3,026 337 685
S3 3,033 337 678
S4 3,049 339 660
S5 3,029 337 682
S6 3,049 339 660
Table 21: Data Count for Various Data Splitting Strategies.
Appendix F Data Splitting Strategies and Statistics
Table 21 presents the data counts for the time-series split and leave-one-out methods for different students in the train, validation and test sets.
Appendix G Distance Metric
The exact match metric is overly strict, so we also consider cases where the model’s predicted errors slightly deviate from the annotated errors. For example, minor character differences, though not entirely correct, may still be acceptable. Similarly, the model might predict the step immediately before or after the actual error. The distance metric is designed to distinguish these minor deviations from predictions that are entirely misaligned with the annotated errors, providing a more comprehensive evaluation of the model’s ability to detect errors accurately.
To evaluate the discrepancy between the predicted erroneous equations $P_{i}$ and the annotated erroneous equations $G_{i}$, we compute the distance $D(P_{i},G_{i})$ for each pair. The positions of these equations within the student’s solution are represented by $P_{\text{start}}$ and $P_{\text{end}}$ for $P_{i}$, and $G_{\text{start}}$ and $G_{\text{end}}$ for $G_{i}$. Specifically, $P_{\text{start}}$ is the starting character position of $P_{i}$, and $P_{\text{end}}$ is the ending position. Similarly, $G_{\text{start}}$ is the starting position of $G_{i}$, and $G_{\text{end}}$ is the ending position.
A distance of 0 is assigned when the model accurately identifies the error. If the model incorrectly classifies an incorrect answer as correct, a penalty of 127 is applied, representing the average length of the student’s solution. Specifically, the calculation of $D(P_{i},G_{i})$ is defined as follows:
  1. 1.
    If both $P_{i}$ and $G_{i}$ are “None” (indicating no errors in the equation), the distance $D(P_{i},G_{i})=0$.
  2. 2.
    If only one of $P_{i}$ or $G_{i}$ is “None,” the distance is set to the penalty value $D_{\text{penalty}}$.
  3. 3.
    If the start and end positions of $P_{i}$ and $G_{i}$ perfectly match, the distance $D(P_{i},G_{i})=0$.
  4. 4.
    When $P_{i}$ and $G_{i}$ do not overlap, the distance is the character difference between their boundaries:
    • If $P_{\text{end}}\leq G_{\text{start}}$, then $D(P_{i},G_{i})=G_{\text{start}}-P_{\text{end}}$.
    • If $G_{\text{end}}\leq P_{\text{start}}$, then $D(P_{i},G_{i})=P_{\text{start}}-G_{\text{end}}$.
  5. 5.
    When $P_{i}$ and $G_{i}$ partially overlap, the distance is the sum of the non-overlapping character differences:
    • If $G_{\text{start}}\leq P_{\text{start}}\leq P_{\text{end}}\leq G_{\text{end}}$, then $D(P_{i},G_{i})=(P_{\text{start}}-G_{\text{start}})+(G_{\text{end}}-P_{\text{end}})$.
    • If $P_{\text{start}}\leq G_{\text{start}}\leq G_{\text{end}}\leq P_{\text{end}}$, then $D(P_{i},G_{i})=(G_{\text{start}}-P_{\text{start}})+(P_{\text{end}}-G_{\text{end}})$.
The distance is calculated for each pair of $P_{i}$ and $G_{i}$, and the minimum distance is taken as the distance for the corresponding equation pair. Formally, the distance is measured as follow:
$D(P_{i},G_{i})=$ (2)
$\begin{cases}0,&\text{if both }P_{i}\text{ and }G_{i}\text{ are ``None''}\\ \text{min}(D_{\text{penalty}},\text{ret\_dis}),&\text{if one of }P_{i}\text{ or }G_{i}\text{ is ``None''}\\ 0,&\text{if }P_{i}\text{ and }G_{i}\text{ perfectly match}\\ G_{\text{start}}-P_{\text{end}},&\text{if }P_{\text{end}}\leq G_{\text{start}}\\ P_{\text{start}}-G_{\text{end}},&\text{if }G_{\text{end}}\leq P_{\text{start}}\\ (P_{\text{start}}-G_{\text{start}})+(G_{\text{end}}-P_{\text{end}}),&\text{if }G_{\text{start}}\leq P_{\text{start}}\leq P_{\text{end}}\leq G_{\text{end}}\\ (G_{\text{start}}-P_{\text{start}})+(P_{\text{end}}-G_{\text{end}}),&\text{if }P_{\text{start}}\leq G_{\text{start}}\leq G_{\text{end}}\leq P_{\text{end}}\\ \text{min}(D_{\text{penalty}},\text{ret\_dis}),&\text{otherwise.}\end{cases}$
$\text{Distance}=\frac{1}{n}\sum_{i=1}^{n}D(P_{i},G_{i})$ (3)
where $n$ is the number of instance.
Appendix H o1-mini Result
The o1-mini model, developed by OpenAI, is renowned for its strong mathematical reasoning capabilities. To evaluate its performance, we conducted experiments using a time series split setting, where inference was performed using few-shot prompts without rationales. We analyzed its performance across three tasks and compared the results with our fine-tuned end-to-end model.
The results of the answer accuracy assessment task are presented in Table 22. Notably, o1-mini demonstrates exceptional mathematical capabilities, achieving an overall accuracy of 92.54% in identifying the correctness of the student’s answers. This surpasses the best-performing end-to-end model with rationale.
The results of the error identification task are shown in Table 23. o1-mini demonstrated stronger mathematical reasoning and understanding abilities, achieving a 41.98% exact match ratio, surpassing the end-to-end model. However, it performed worse in terms of the distance metric. Upon reviewing o1-mini’s outputs, we found that in cases where errors were due to unfinished problems, the student’s calculations were correct. Teachers typically did not mark any formulas as incorrect. Despite this, o1-mini still flagged incorrect formulas in these cases, which introduced significant penalties in the distance metric, leading to a poorer performance in this aspect.
The results of the feedback generation task are shown in Table 24. R-1, R-2, R-L, and Rating represent ROUGE-1, ROUGE-2, ROUGE-L, and LLM Rating metrics, respectively. o1-mini outperformed the end-to-end model in terms of BERTScore. However, its responses were often overly lengthy, generating additional irrelevant information, which impacted the evaluation results, and limited it performance compared to GPT-3.5. In the LLM rating conducted by using GPT-4, o1-mini received an almost perfect score of 4.70. By contrast, in the evaluation conducted by human annotators (i.e., Table 8), o1-mini’s performance was less satisfactory in generating adaptive feedback, as its responses were often overly lengthy and lacked focus, reducing their effectiveness in addressing specific student errors.
All General Gain Physics Geo. Prob. Other
o1-mini w/o $r$ 92.54% 92.24% 92.79% 92.40% 90.00% 93.75% 96.97%
End-to-end w/o $r$ 87.73% 84.05% 93.69% 90.06% 90.00% 62.50% 90.91%
End-to-end w/ $r$ 91.54% 88.79% 95.50% 93.57% 95.00% 62.50% 96.97%
Table 22: o1-mini’s Results of Answer Accuracy Assessment.
EM ($\uparrow$) Distance ($\downarrow$)
o1-mini w/o $r$ 41.98% 56.30
End-to-end w/o $r$ 30.25% 71.48
End-to-end w/ $r$ 40.12% 51.78
Table 23: o1-mini’s Results of Problem-Solving Error Identification.
R-1 R-2 R-L BERTScore Rating
o1-mini w/o $r$ 0.1956 0.0633 0.1839 0.7986 4.7
End-to-end w/o $r$ 0.1862 0.0751 0.1761 0.5327 0.99
End-to-end w/ $r$ 0.2842 0.1418 0.2703 0.6786 1.91
Table 24: o1-mini’s Results of feedback generation.
Appendix I Mathematical Problem-Solving Ability of the Model
To evaluate the problem-solving capabilities of the models used in the baseline with respect to the questions in our dataset, we employed the test set from the time series split method. The results, presented in Table 25, show that even the smallest parameter model, Llama 8B, achieved an overall accuracy of 75.95%. In contrast, Llama 80B attained an impressive overall accuracy of 92.37%, surpassing all of our students. This demonstrates that the LLMs we utilized are capable of understanding and solving the problems present in this task.
Category Questions LLaMA3 8B LLaMA3 70B GPT-3.5
All 603 75.95% 92.37% 82.59%
General 232 75.43% 90.51% 84.91%
Gain 111 80.18% 93.69% 88.29%
Physics 171 73.68% 95.32% 78.36%
Geometry 40 77.50% 87.50% 77.50%
Probability 16 75.00% 100.00% 81.25%
Other 33 75.76% 87.88% 75.76%
Table 25: Performance of Different Models Across Various Problem Categories.
S1 S2 S3 S4 S5 S6
Llama3 8B w/o $r$ 59.59% 70.36% 61.50% 67.88% 66.72% 55.15%
w/ $r$ 62.37% 72.26% 64.90% 71.06% 73.90% 66.97%
Llama3 70B w/o $r$ 84.77% 88.91% 84.96% 82.88% 87.24% 82.73%
w/ $r$ 87.70% 86.13% 83.78% 81.67% 81.67% 83.79%
GPT-3.5 w/o $r$ 40.85% 47.45% 45.28% 36.36% 51.32% 28.94%
w/ $r$ 50.07% 65.99% 50.74% 47.12% 52.05% 29.55%
Single-task w/o $r$ 20.06% 87.88% 85.25% 67.73% 85.19% 85.45%
w/ $r$ 93.85% 94.01% 92.04% 89.85% 92.08% 92.88%
Multi-task w/o $r$ 0.00% 30.80% 66.08% 70.61% 90.91% 0.00%
w/ $r$ 89.75% 88.03% 47.20% 0.00% 85.34% 18.33%
End-to-end w/o $r$ 87.12% 57.96% 85.55% 87.73% 90.03% 89.09%
w/ $r$ 94.58% 95.33% 90.86% 89.85% 96.48% 91.67%
Table 26: Model Performance on Answer Accuracy Assessment Across Different Students.
Appendix J Impact of Variations in Student Problem-Solving Styles on Model Grading
The problem-solving results in our dataset exhibit considerable diversity among students, as detailed in Table 3. Students 2, 3, and 4 provided more detailed and complex solutions, often including natural language explanations. In contrast, Students 1, 5, and 6 used fewer equations and offered more abbreviated responses, with Students 1 and 6 almost entirely omitting natural language descriptions of their processes. The variation in student response styles raises the question of whether this diversity impacts the models’ grading performance.
We can examine this question by looking at the experimental results in Table 26 under the leave-one-out data split method. For the non-finetuned models, the number of parameters plays a significant role in their ability to interpret and assess students’ problem-solving results. The Llama3 8B model, with fewer parameters, struggled to accurately grade the concise and less descriptive responses provided by Students 1 and 6, resulting in a lower performance. Conversely, the Llama3 70B model, although still affected by these abbreviated responses, demonstrated greater resilience due to its enhanced comprehension capabilities. Similarly, GPT-3.5, despite lower overall accuracy due to its more stringent evaluation of student processes, exhibited comparable trends across the different students, aligning with Llama3 in terms of accuracy variation.
For the fine-tuned models, those trained with the multi-task learning method demonstrated less robustness. We found that these models occasionally failed to provide any output when confronted with particularly brief responses, leading to instances of zero accuracy. Moreover, the fine-tuned models showed reduced performance when assessing more detailed and complex solutions, such as those from Student 2, likely due to limitations in their ability to process and understand intricate reasoning. However, when models were allowed to reference rationales during answer accuracy assessment, their overall performance improved, and discrepancies in grading accuracy across different problem-solving styles were mitigated. That means the incorporation of rationales can assist models in navigating the challenges posed by varied student responses, enabling them to more accurately determine whether a student’s problem-solving process is correct.
To sum up, the experiment indicates that the diversity in students’ problem-solving styles affect the models’ answer accuracy assessment performance. If the student’s solution includes more detailed explanations or even natural language descriptions, the LLMs are better able to evaluate the solution accurately, leading to more reliable grading outcomes.
Appendix K Evaluation Criteria for Generated Feedback Quality

### v2

@@ 8 Conclusion
This study focuses on the application of generative AI in mathematics education, particularly its reliability in grading authentic student problem-solving processes and generating feedback. To facilitate a rigorous analysis, we construct the MathEDU dataset, which includes real student problem-solving processes and teacher feedback for GRE-level math problems. A range of models is evaluated on answer correctness classification, error identification, and feedback generation. Experimental results show that our fine-tuned models achieve notable improvements in classifying correctness and identifying erroneous steps, while LRMs demonstrate stronger overall performance across the three tasks. Nevertheless, current models still struggle with generating personalized suggestions. Common issues include imprecise wording, numerical miscalculations, omission of critical intermediate steps, and lengthy yet untargeted responses. An advanced method is required to build pedagogy-aware AI systems that can both produce reliable, adaptive guidance and better interpret students’ mathematical reasoning, which we leave as future work. Additionally, extending the scale and scope to a broader range of problem types and domains is left as our future work.
@@
Limited Model and Method Exploration. Our study focused on Llama3 8B, Llama3 70B, and GPT-3.5, with experiments limited to few-shot prompting and LoRA-based fine-tuning. More recent LLMs remain unexplored, as keeping pace with rapidly evolving releases and conducting reliable expert-based evaluations poses substantial challenges. To ensure consistency and completeness, we focused on a representative set of models in this study. Additionally, our fine-tuning methods did not delve into incorporating students’ individual learning progress, such as tracking their mastered concepts. Despite these limitations, we hope our work can contribute to the growing exploration of LLM applications in mathematics education and encourage further research to refine and expand these methods.
@@
To address potential privacy concerns associated with the MathEDU dataset, this section outlines the measures taken to protect annotators’ privacy and the ethical considerations in releasing the dataset. The MathEDU dataset contains problem-solving records annotated by students from diverse academic backgrounds. However, no personally identifiable information, such as names or IDs, was collected. Each record is assigned an anonymized identifier to categorize entries from the same student. While the dataset includes the students’ academic backgrounds (i.e., majors), this information is provided solely for analytical purposes. As highlighted in previous sections, understanding how problem-solving styles vary across mathematical abilities and expertise is a valuable area of research. All annotators involved in the task, including students and mathematics experts serving as teachers, were compensated based on the minimum hourly wage standards of their respective countries. The student annotators were informed that their personal data would remain confidential and that only their academic background would be disclosed. This consent was obtained prior to participation in the annotation task. The dataset’s structure ensures that individual privacy is preserved while facilitating research in educational applications. In alignment with ethical guidelines, the dataset will only be released for research purposes.
Acknowledgments
This research was partially supported by National Science and Technology Council, Taiwan, under grant NSTC 114-2221-E-A49-057-MY3.
@@ Appendix A Math Word Problem Dataset Selection
Goal:
The goal is to obtain students’ problem-solving processes when they make errors in solving math problems, which will help train language models for applications in math education.
Annotation Process:
A Google Sheet link will be provided to the students, which will contain both the Chinese and English versions of the questions. Students can choose one of the following methods to complete their solutions:
1: Solve on a tablet. 2: Solve on paper and take a picture afterward (one picture per question). After completing the solutions, students should name the file as <problem_id> (e.g., 2456.jpg) and upload it to the cloud drive.
The timeline for completing 750 questions will be discussed with students to set a deadline (estimated to be within two months). Weekly meetings will be scheduled to report progress and understand the students’ problem-solving status. There will be no weekly minimum requirement; students can allocate their time as they see fit.
Annotation Guidelines:
All problems are junior high school-level math problems. Students must show the full solution process and the final answer with the following requirements:
- Clearly write the solution (top-down).
- Include simple English explanations, no Chinese.
- Example 1 :
[…]
- Example 2 :
[…]
- Round to 2 decimal places. Use 3.14 for pi. Simplify fractions. No calculators allowed.
- If you encounter a question you cannot solve, note the problem number and the type of error, then skip it. Error types include:
1: Don’t know how to solve it. 2: Problem definition is unclear. 3: Unrecognizable symbols or notation. 4: Problem requires a diagram, but none is provided. 5: Other…
- If you feel the instructions are unclear or have any other questions, please ask the responsible staff immediately.
Table 9: Math Problem-Solving Process Guidelines.
@@ B.2 Academic Backgrounds and Question Distribution of Student Annotators
To collect data on math problem-solving, we invited six university students to participate in the annotation process. Table 10 presents the majors of the student annotators and the distribution of questions they answered. We aimed for diversity by including individuals from different academic backgrounds, such as Japanese Studies and Applied Mathematics, to capture a range of math proficiency levels in the problem-solving process. Before the annotation process, students were asked to attempt 10 questions to ensure their responses adhered to the required format. This step helped avoid issues such as incomplete equations or disorganized layouts, facilitating smoother data collection.
ID Major Total General Gain Physics Geometry Probability Other
S1 Applied Mathematics 683 276 (40.41%) 132 (19.33%) 182 (26.64%) 49 (7.17%) 7 (1.02%) 37 (5.42%)
S2 Finance 685 297 (43.36%) 133 (19.42%) 188 (27.45%) 39 (5.69%) 4 (0.58%) 24 (3.50%)
S3 Japanese 678 281 (41.45%) 136 (20.06%) 180 (26.55%) 42 (6.19%) 7 (1.03%) 32 (4.72%)
S4 Information Management 660 268 (40.60%) 140 (21.21%) 167 (25.30%) 46 (6.97%) 8 (1.21%) 31 (4.70%)
S5 Mathematics Education 682 275 (40.32%) 139 (20.38%) 186 (27.27%) 41 (6.01%) 8 (1.17%) 33 (4.84%)
S6 Physics 660 262 (39.70%) 127 (19.24%) 175 (26.52%) 55 (8.33%) 7 (1.06%) 34 (5.15%)
Table 10: Student Majors and Number of Questions Answered.
@@ Appendix C Supplementary Information for Each Error Category
The detailed definitions of these five error categories are as follows: (1) Wrong Mathematical Operation/Concept refers to the use of incorrect operations or inappropriate concepts. (2) Calculation Error involves mistakes in arithmetic, equation solving, or unit conversion. (3) Incomplete Answer occurs when a student begins a correct procedure but does not finish it. (4) Careless Error captures inattentive mistakes, such as incorrect number substitution or missing digits. (5) Lack of Necessary Mathematical Concepts indicates errors stemming from insufficient foundational knowledge. While both (1) and (5) involve conceptual issues, the former reflects misapplication, whereas the latter involves an inability to begin due to missing knowledge. The following are examples of each error category.
@@
[…]
@@
Appendix D Student Problem-Solving Records
Table 16 summarizes the correctness of students’ solutions together with their problem-solving behaviors (e.g., equations and words per problem), illustrating the diversity of data collected even from only six participants.
Appendix E Additional Results on Feedback Generation
Table 17 reports supplementary results for the feedback generation task using ROUGE-1 and ROUGE-2. These metrics reflect n-gram level lexical overlap, which may be less aligned with semantic fidelity and pedagogical usefulness. We therefore include them here for completeness, while focusing on ROUGE-L, BERTScore, and LLM Rating in the main text.
R-1 R-2 R-L
Llama3 8B w/o $r$ 0.2593 0.0809 0.2391
w/ $r$ 0.1819 0.0557 0.1632
Llama3 70B w/o $r$ 0.2587 0.0881 0.2439
w/ $r$ 0.2456 0.0876 0.2311
GPT-3.5 w/o $r$ 0.2897 0.1041 0.2695
w/ $r$ 0.2848 0.1025 0.2597
o1-mini w/o $r$ 0.1956 0.0633 0.1839
Single-task w/o $r$ 0.1661 0.0730 0.1594
w/ $r$ 0.0790 0.0378 0.0750
Multi-task w/o $r$ 0.0130 0.0049 0.0124
w/ $r$ 0.1650 0.0719 0.1552
End-to-end w/o $r$ 0.1862 0.0751 0.1761
w/ $r$ 0.2842 0.1418 0.2703
Table 17: ROUGE results (R-1, R-2, R-L) for the feedback generation task.
Appendix F Evaluation Criteria for Generated Feedback Quality